Chris bought 1 and 3/4 pounds of nuts. Two-thirds of this amount was used at a dinner party. How many pounds of nuts were used?
Answer by mananth(16946)

1 (3/4) pounds
2/3 was used
..
1 3/4 = 7/4
..
2/3 * 7/4 = 7/6 was used.
1 1/6 pound was used.